OpenStudy (anonymous):

Solve x − 2y = 6 3x − 6y = 18

OpenStudy (anonymous):

(2, −2) (3, negative three halves) No solutions Infinitely many solutions

OpenStudy (anonymous):

@iGreen

OpenStudy (eric_d):

x-2y=6 x=6+2y...............1 3x-6y=18 x=6+2y............2 Substitute 1 into 2 to find value of x first

OpenStudy (eric_d):

sorry, nt x but to find value of y

OpenStudy (eric_d):

what will u get @_angeliquexoxo

OpenStudy (anonymous):

I am actually confused... how do I find the value of y?

OpenStudy (jumperman):

Well you need to get y by itself right?

OpenStudy (jumperman):

Don't you use the inverse of operations for that?

OpenStudy (anonymous):

How do I do the inverse of operations?

OpenStudy (jumperman):

That's a great question. Sometimes it helps with just taking some parts of the equation, not all parts, but some and just reversing the sign. Instead of subtracting you add. Other times depending on what type of equation it is, you have to be careful what you change.

OpenStudy (jumperman):

But inverse of operations is Sadmep. Subtract, add, divide, multiply, exponents and the parentheses. It's the opposite of Pemdas.

OpenStudy (anonymous):

Okay so what do I need to do?

OpenStudy (jumperman):

Well, I think, if this is the right type of equation is that x will always equal one, when there is no number attached to it. so I think start there.

OpenStudy (jumperman):

Then try and change the signs after that? It depends ultimately on which ones you need to change.

OpenStudy (eric_d):

6+2y=6+2y 2y-2y=0 0y=0 No solution! am I rite @mathmale

OpenStudy (anonymous):

There had to be a solution

OpenStudy (mathmale):

This whole discussion is about solving a system of linear equations. This means finding a point (x,y) that satisfies BOTH of the given equations. Someone suggested using the method of elimination by substitution. That is why he/she has suggested taking one of the two given equations and solving it for y algebraically. If we take the first given equation, x-2y=6, and "solve for x," here are the steps of that procedure: x-2y=6 Isolate the -2y term by subtracting x from both sides of this equation.

OpenStudy (mathmale):

That results in -2y=6-x We want y alone, not -2y, so we divide both sides of the above equation by -2 to obtain y=(6-x)/(-2)

OpenStudy (mathmale):

We can now substitute (6-x)/(-2) for ' y ' in the second equation given. This second equation is 3x-6y=18. Rewrite this as 3x-6( y )=18. Now you have a formula for y. Substitute the right side of that formula inside the parentheses shown above; that is, substitute (6-x)/(-2) for y in the above equation. Please type out the result.